# Tazza: Shuffling Neural Network Parameters for Secure and Private Federated Learning

### v1

@@ 7.3. Confidentiality Attack Mitigation
Table 1. Average and highest PSNR/LPIPS between original and inferred sample results from confidentiality attack for different dataset/defense configurations.
@@ 7.4. Integrity Attack Mitigation
Figure 10. (a) Global model accuracy and backdoor accuracy for different defense schemes against backdoor attack. (b) LPIPS and accuracy of different defense mechanisms under noise injection attack.
In this section, we evaluate Tazza’s robustness against various integrity attacks, specifically focusing on noise addition and backdoor attack scenarios. The experiments exploit the CIFAR10-Mixer configuration with medium-level DPSGD as the confidentiality defense mechanism for all baselines. For the backdoor attack, attackers assign specific labels to samples by injecting a 4$\times$4 red square patch into the upper-left corner of 50% of their training data as the backdoor trigger (Lee et al., 2024). For the noise addition attack, Gaussian noise with a scaling factor of 0.25 is injected into the model weights before they are uploaded to the server.
@@ 7.5. Scalability on Real Embedded Platforms
Figure 12. Latency of different confidentiality attack mitigation schemes with varying (a) hidden layer counts and (b) number of local training updates.
@@
Figure 13. Latency of different confidentiality attack mitigation schemes with varying model architectures
@@
Finally, Figure 14 depicts the latency of various defense schemes operating on different embedded platforms using an MLP model with a single hidden layer and 10 mini-batch updates. Overall, we can observe that Tazza shows superior performance compared to alternative mitigation approaches, by Tazza achieving up to 6.70$\times$ lower latency compared to gradient pruning on the Raspberry Pi 3B. Furthermore, as a positive side-effect, Figure 14 shows that Tazza effectively reduces computational disparities among heterogeneous client platforms, with a 2.01$\times$ difference in latency between the most powerful and weakest devices, compared to 2.68$\times$ for gradient pruning. These results serve as evidence that Tazza is suitable for mobile/embedded federated learning systems.
@@ 8. Discussion
$\bullet$ Permutation Variant Neural Networks. This work proposes Tazza, a framework addressing two core security challenges in federated learning by leveraging the permutation equivariance and invariance properties of neural networks. While these properties are commonly observed in modern neural networks, certain architectures (e.g., CNN) that lack these properties still persist (Battaglia et al., 2018). While they can often be replaced with alternatives that exhibit these properties, there are cases where the model best suited for a specific application does not qualify this criteria. As a result, these models cannot directly leverage the weight shuffling mechanism in Tazza. However, it is important to note that many of these neural networks are optimized at the software level and are eventually transformed into matrix multiplication operations (Chellapilla et al., 2006; Vasudevan et al., 2017). Considering such optimization, it is feasible to devise new shuffling strategies tailored to these specific models, enabling their integration into the Tazza framework.
$\bullet$ Counter Attack Against Tazza. A key function for Tazza to function effectively is ensuring confidentiality by preventing the server from knowing the shuffling rules. However, the server could attempt counter-attacks to infer this information. For instance, there have been attempts to train models capable of solving jigsaw puzzles to recover original inputs from shuffled data (Li et al., 2021). While such studies have made progress in identifying patch-level shuffling, research in this domain remains incomplete (Carlucci et al., 2019). Furthermore, the server might try to infer shuffling rules through neural network parameter analysis. Nevertheless, as discussed earlier, leveraging the orthogonality of frameworks provides a promising direction to counter such attacks. While an intriguing topic, the primary goal of this study is to demonstrate the feasibility of addressing the two critical security issues in federated learning by exploiting permutation invariance and equivariance. As such, we defer this exploration to future work.

### v2

@@ 7.3. Confidentiality Attack Mitigation
Table 2. Average and best PSNR/LPIPS between original and inferred sample results from confidentiality attack for different dataset/defense configurations.
@@ 7.4. Integrity Attack Mitigation
Figure 12. (a) Backdoor accuracy and global model accuracy for different defense schemes against backdoor attack. (b) LPIPS and accuracy of different defense mechanisms under noise injection attack.
In this section, we evaluate Tazza’s robustness against various integrity attacks, specifically focusing on noise addition and backdoor attack scenarios. The experiments exploit the CIFAR10-Mixer configuration with medium-level DPSGD as the confidentiality defense mechanism for all baselines. For the backdoor attack, attackers assign specific labels to samples by injecting a 4$\times$4 red square patch into the upper-left corner of 50% of their training data as the backdoor trigger (Lee et al., 2024b). For the noise addition attack, Gaussian noise with a scaling factor of 0.25 is injected into the model weights before they are uploaded to the server.
@@ 7.5. Scalability on Real Embedded Platforms
Figure 14. Latency of different confidentiality attack mitigation schemes with varying (a) hidden layer counts and (b) the number of local training updates.
@@
Figure 15. Latency of different confidentiality attack mitigation schemes with varying model architectures.
@@
Figure 16 depicts the latency of various defense schemes operating on different embedded platforms using an MLP model with a single hidden layer and 10 mini-batch updates. Overall, we can observe that Tazza shows superior performance compared to alternative mitigation approaches, by Tazza achieving up to 6.70$\times$ lower latency compared to gradient pruning on the Raspberry Pi 3B. Furthermore, as a positive side-effect, Figure 16 shows that Tazza effectively reduces computational disparities among heterogeneous client platforms, with a 2.01$\times$ difference in latency between the most powerful and weakest devices, compared to 2.68$\times$ for gradient pruning. These results serve as evidence that Tazza is suitable for mobile/embedded federated learning systems.
7.6. Impact of Adaptive Attacks
In this section, we demonstrate the robustness of Tazza against potential adaptive attacks, where an attacker possesses knowledge of the system and attempts to exploit it to bypass the defense mechanism.
7.6.1. Parameter-based shuffling rule inference
Figure 17. (a) Average Error (L1-norm) between server inferred and the authentic shuffling order. (b) Confidentiality attack results showing the bottom-5 LPIPS values with and without the inference of the shuffling rule.
A key point to assure that Tazza operates as intended is to ensure the confidentiality of the shuffling rules from the server. However, it is practical to assume that the server would attempt counterattacks to infer this information. For instance, the server might try to deduce shuffling rules through neural network parameter analysis. As an initial study, we evaluate Tazza’s robustness against such attacks using an advanced adversary employing additional shuffling rule inference techniques. In this scenario, the attacker tries to identify a shuffling rule that minimizes the difference of the parameters (i.e., L2-norm) between the shuffled model parameters from the client against the global model.
To examine this, we used the MNIST dataset with 3-layered MLP models trained over a single update. Shuffling was applied only to the input layer, while the order of other layers remained unchanged, representing the simplest and most vulnerable configuration within our setup. Figure 17 (a) plots the distribution of average errors (L1-norm) between the inferred and ground truth shuffling rules across 100 clients. Note that we use the L1-norm since the shuffling rule can be interpreted as the indices of a given matrix. As the plots show, the average error was concentrated near 280, indicating that the index of a given row was located 280 rows off on average. This demonstrates the robustness of Tazza against a shuffling rule inference attack issued at the server. Figure 17 (b) visualizes the original samples that exhibited bottom-5 LPIPS values (5 most similar attack results) and the inferred samples via an inverting gradient attack with and without the shuffling rule inference. As the figure shows, the inferred samples are semantically uninterpretable, demonstrating the robustness of Tazza to adaptive attacks that try to infer the shuffling rule based on model parameters.
Figure 18. Results of the jigsaw puzzle solver-based image restoration attack.
7.6.2. Jigsaw puzzle-based shuffling rule inference
In addition to rule inference using the parameters, there have also been efforts to train models capable of solving jigsaw puzzles to reconstruct original inputs from shuffled data (Li et al., 2021). While these studies have shown success in recovering shuffled data with a relatively small number of patches (e.g., $3\times 3$), research on handling a large number of patches commonly used in neural networks (e.g., $14\times 14$, $16\times 16$, $32\times 32$) and pixel-level shuffling remains mostly unexplored (Carlucci et al., 2019).
To evaluate the robustness of Tazza against such adaptive attacks, we conduct stress tests using the STL dataset and a transformer-based jigsaw puzzle solver (Chen et al., 2023). Since these models are designed for high-resolution images ($>350\times 350$) with a small number of patches (e.g., $3\times 3$), we adopt the same configuration for this experiment. Figure 18 presents the results of image restoration attacks using the jigsaw puzzle solver. As the figure shows, the attack model successfully reconstructs certain aspects of the original images (‘Patch Shuffled Result’ in Figure 18). However, when both patch- and pixel-level shuffling are applied simultaneously, the model fails to recover the correct patch locations (‘Patch+Pixel Shuffled Result’ in Figure 18). Moreover, the semantic information becomes unrecognizable due to intra-patch pixel shuffling, which indicates Tazza’s resilience against such attacks.
Furthermore, even under relatively simple configurations, such as the MNIST dataset, the number of possible shuffling rules grows factorially (e.g., $(28\times 28)!>10^{1930}$), making it impractical for an attacker to recover the correct rule. For real-world datasets with high dimensional inputs, the search space of the shuffling rule expands exponentially, significantly increasing both the computational burden and infeasibility of launching a successful attack.
@@ 8. Discussion
$\bullet$ Permutation Variant Neural Networks. This work proposes Tazza, a framework addressing two core security challenges in federated learning by leveraging the permutation equivariance and invariance properties of neural networks. While these properties are commonly observed in modern neural networks, certain architectures (e.g., CNN) that lack these properties still persist (Battaglia et al., 2018). While they can often be replaced with alternatives that exhibit these properties, there are cases where the model best suited for a specific application does not qualify these criteria. As a result, these models cannot directly leverage the weight shuffling mechanism in Tazza. However, it is important to note that many of these neural networks are optimized at the software level and are eventually transformed into matrix multiplication operations (Chellapilla et al., 2006; Vasudevan et al., 2017). Considering such optimization, it is feasible to devise new shuffling strategies tailored to these specific models, enabling their integration into the Tazza framework.
$\bullet$ Adaptive Attacks Against Tazza. While we conducted an evaluation on two types of adaptive attacks, there remain potential adaptive strategies that malicious attackers could employ. For instance, attackers might use tampered shuffling rules before uploading their model to the server or leak the shuffling rule to the server. In the case of attackers exploiting tampered shuffling rules, the tampered model is unlikely to impact benign clients, as the clustering algorithms employed by the server can effectively filter out such malicious models. Additionally, the server can enhance the integrity of the federated learning pipeline by incorporating a zero-knowledge proof (ZKP) approach to verify that the shuffled weights have not been tampered with.
In the (unlikely) case where the shuffling rule is leaked, clients can implement additional security measures, such as secure multi-party computation (SMPC), during the rule-sharing process. This ensures that each client holds only partial information about the shuffling rule, thereby preventing attackers from gaining full knowledge of the rule without the cooperation of benign clients. While these approaches are promising and effective, they introduce additional computational and communication overhead. Consequently, designing efficient and effective security protocols or features represents an important and compelling direction for future research.